find a series solution in powers of x of the equation

2x2d2y/dx2+xdy/dx+(x2-1)y=0


1
Expert's answer
2021-07-05T14:21:54-0400

Solution

Let

"y\left(x\right)=\sum_{n=0}^{\infty}{a_nx^n}"

"\frac{dy}{dx}=\sum_{n=0}^{\infty}{\left(n+1\right)a_{n+1}x^n}"

"\frac{d^2y}{dx^2}=\sum_{n=0}^{\infty}{\left(n+1\right)\left(n+2\right)a_{n+2}x^n}"

Substitution into equation:

"2\sum_{n=0}^{\infty}{\left(n+1\right)\left(n+2\right)a_{n+2}x^{n+2}}+\sum_{n=0}^{\infty}{\left(n+1\right)a_{n+1}x^{n+1}}+\sum_{n=0}^{\infty}{a_nx^{n+2}}-\sum_{n=0}^{\infty}{a_nx^n}=0"

"a_1x-a_0-a_1x+2\sum_{n=2}^{\infty}{n\left(n-1\right)a_nx^n}+\sum_{n=2}^{\infty}{na_nx^n}+\sum_{n=2}^{\infty}{a_{n-2}x^n}-\sum_{n=2}^{\infty}{a_nx^n}=0"

Coefficient near xn are equal to zero.

n=0: a0=0

n=1: 0=0 for any a1  

n>1: 2n(n-1)an+nan+an-2-an=0 => (2n2-n-1)an+ an-2 = 0 =>

an =-an-2/ (2n2-n-1)= -an-2/ (2n2-n-1) = -an-2/ [(2n+1)(n-1)]

So a2k=0 for any k≥0

and finally a2k+1  = -a2k-1/ [(4k+3)2k] for k>0

Solution is

"y(x)=\sum_{k=1}^{\infty}{a_{2k-1}x^{2k-1}}"